# ModelShield: Adaptive and Robust Watermark against Model Extraction Attack

### v2

@@ IV-H1 Editing Attack
We can initially confirm that attacking watermarked data will undoubtedly impact the performance of watermark detection. However, as observed from the Fig 8, even after various editing operations on watermarked data, the imitation models trained using our watermarking method still show a statistically significant difference in sentence watermark score compared to $M_{L}$ and $M_{O}$. Moreover, our watermarking method does not compromise the original task performance of the model despite being subjected to editing attacks. The task performance of the imitation model remains comparably high, akin to scenarios without such attacks.
IV-H2 Selective Training on Watermarked Data Subsets
(a) SWS Distribution for Llama2 on HC3
(b) SWS Distribution for Llama2 on WILD
Figure 10: Attackers combine non-watermarked data from other sources for training imitation models: In this scenario, attackers use all the watermark data obtained from queries and mix in non-watermarked data from other sources to train imitation models. We controlled the watermark data at a fixed value (400 entries) and mixed in non-watermarked data, varying the watermark data’s proportion $\rho$ to 10%, 25%, 50% and 75%.
In this mixed training scenario, we prepared six distinct datasets, each containing a varying ratio $\eta$. The values of $\eta$ were set at 0, 0.1, 0.25, 0.5, and 0.75, and we ensured that the size of the total watermark dataset $|A|$, consistently remained at 4000 instances. The outputs of the suspect model were then evaluated based on our comprehensive knowledge of the query history.
@@ IV-H2 Selective Training on Watermarked Data Subsets
Even though attackers use only a portion of the watermarked data for training, we conduct watermark detection on all the queries $Q$ known to have been scraped by the victim model’s owner. As depicted in Fig 9, a distinct statistical discrepancy is apparent between models trained with and without watermarked data (with a p-value of about $10^{-2}$), even when only 10% of the watermark data (400 samples) are used for imitation model training.
@@
This not only demonstrates the robustness of our watermark verification method but also indicates the learnable nature of the watermarked data. It can also be seen that as $\eta$ increases, the average sentence watermark score rises accordingly. This aligns with our intuition that the more watermarked data used for training the imitation model, the easier it is to perform watermark detection.
IV-H3 Mixture of unwatermarked data
@@
In a scenario similar to selective training on watermarked data subsets, attackers might mix all queried data from the victim model with additional unwatermarked source data for training the imitation model. We set $A^{D}=A$ in this case. Our experiments maintain a fixed number of watermarked data while varying the total training data $|D|$, altering $\rho$ to values of 0.1, 0.25, 0.5 and 0.75. We only detect watermarks on queries known to the victim model’s owner since the attacker can blend in any unwatermarked data. The results, as illustrated in Fig 10, reveal that increasing the proportion of watermarked data correspondingly elevates the watermark detection in the text generated by the imitation model. Watermarks are detectable in all cases with p-values below $10^{-2}$, without affecting the performance on downstream QA tasks.
IV-I LLM Generation Visualization
@@
As illustrated in Fig 11, we analyze the BertScore [63] comparisons between the imitation model’s outputs and the original foundational model’s responses to identical queries. It’s evident that the output of the imitation model, in various text relevance scenarios, consistently shows high sentence watermark scores, proving that the watermark can effectively transfer to the imitation model and be effectively detected, making it a necessary mechanism. Conversely, outputs from base models (represented in blue) show generally low sentence watermark scores, proving that the watermark is not prone to false positives.
@@ V Conclusion
In this paper, we propose ModelShield, a plug-and-play watermarking method for language models designed to combat model extraction attacks. Unlike existing IP protection watermarking methods that modify model’s output or distribution with heuristic rules, ModelShield uses system prompt guidance to minimize distortion seamlessly. Leveraging the self-reminding capabilities of LLMs, ModelShield integrates user queries with a self-generated watermark system prompt, subtly embedding watermark words into outputs without altering the model’s internal probabilities or outputs. This approach maintains text quality while ensuring watermark effectiveness and robustness. Extensive experiments validate ModelShield’s efficacy, showing remarkable detection sensitivity even under interference. Its unobtrusiveness and model-agnostic nature make it practical for real-world applications, reducing detection and reverse engineering risks. Although our method can resist some model extraction attacks, it has limitations. Attackers might use jailbreak prompts to bypass preset system instructions, making it uncertain if our watermarking will be compromised. Defending against jailbreak prompts is complex and extensively studied, so we do not delve into whether system instructions will be executed. We hope ModelShield offers fresh insights for copyright protection of large language models. It provides a viable and efficient strategy for model owners to safeguard their IP rights at a low cost, fostering fairness and innovation in the LMaaS market.

### v4

@@ IV-I1 Editing Attack
We can initially confirm that attacking watermarked data will undoubtedly impact the performance of watermark detection. However, as observed from Fig 8, even after various editing operations on watermarked data, the imitation models trained using our watermarking method still show a statistically significant difference in sentence watermark score compared to $M_{L}$ and $M_{O}$ (with detailed detection $p$-values remained consistently below 1e-18). Moreover, our watermarking method does not compromise the original task performance of the model despite being subjected to editing attacks. The task performance of the imitation model remains comparably high, akin to scenarios without such attacks.
IV-I2 Selective Training on Watermarked Data Subsets
(a) WS for Llama2 on HC3
(b) WS for Llama2 on WILD
Figure 10: Attackers use all the watermark data obtained from queries and mix in non-watermarked data from other sources to train imitation models. Using a fixed watermark data of 400 entries, we varied the watermark proportion ($\rho$) to 10%, 25%, 50%, and 75% by mixing in non-watermarked data.
We prepared seven distinct datasets, each containing a varying ratio $\eta$. The values of $\eta$ were set at 0, 0.01, 0.07, 0.1, 0.25, 0.5, and 0.75, and we ensured the size of the total watermark dataset $|A|$, consistently remained at 4000 instances. The outputs of the suspect model were then evaluated based on our comprehensive knowledge of the query history.
@@ IV-I2 Selective Training on Watermarked Data Subsets
Even though attackers use only a portion of the watermarked data for training, we conduct watermark detection on all the queries $Q$ known to have been scraped by the victim model’s owner. As depicted in Fig 9, a distinct statistical discrepancy is apparent between models trained with and without watermarked data (with a $p$-value of about $10^{-2}$ in rapid detection), even when only 10% of the watermark data (400 samples) are used for imitation model training.
@@
This not only demonstrates the robustness of our watermark verification method but also indicates the learnable nature of the watermarked data. It can also be seen that as $\eta$ increases, the average sentence watermark score rises accordingly. This aligns with our intuition that the more watermarked data used for training the imitation model, the easier it is to perform IP infringement detection.
IV-I3 Mixture of unwatermarked data
@@
In a scenario similar to selective training on watermarked data subsets, attackers might mix all queried data from the victim model with additional unwatermarked source data for training the imitation model. We set $A^{D}=A$ in this case. Our experiments maintain a fixed number of watermarked data while varying the total training data $|D|$, altering $\rho$ to values of 0.1, 0.25, 0.5 and 0.75. We only detect watermarks on queries known to the victim model’s owner since the attacker can blend in any unwatermarked data. The results, as illustrated in Fig 10, reveal that increasing the proportion of watermarked data correspondingly elevates the watermark detection in the text generated by the imitation model. Watermarks are consistently detectable with $p$-values below 0.05 in rapid detection and below 1e-8 in detailed detection, without affecting the performance of downstream QA tasks.
TABLE XII: Sensitivity analysis of the detection mechanism under interference: In various interference scenarios and with different thresholds for rapid detection, we report the $p$-values for both rapid detection and detailed detection
Interfered scenarios Rapid Verification Detailed Verification
$\theta=0.07$ $\theta=0.08$ $\theta=0.09$ $\theta=0.10$ $\theta=0.11$ $\theta=0.12$ $\theta=0.13$ comp. $M_{O}$ comp. $M_{L}$
w/o adversarial attack 2.6561e-15 5.9761e-15 1.3617e-14 7.3486e-14 1.7407e-13 4.1777e-13 1.0158e-12 5.7214e-13 9.8524e-43
Random Scrambling 8.8251e-15 7.3274e-14 14.0547e-14 1.9975e-13 1.0355e-12 5.6563e-12 3.2587e-11 7.8541e-21 2.0132e-19
Random Insertion 1.2602e-24 3.4596e-23 1.1082e-22 3.6666e-22 1.2552e-21 4.4538e-21 1.4408e-20 7.6061e-31 2.0132e-23
Random Deletion 7.6379e-24 5.9615e-23 1.7305e-22 5.1614e-22 1.5835e-21 5.0044e-31 1.6311e-20 3.4004e-32 8.5137e-25
Random Replacement 3.5545e-24 3.7567e-23 1.2849e-22 4.5575e-22 1.6793e-21 6.4416e-21 2.5777e-20 5.7236e-31 6.9446e-28
Selective Training 10% 7.7641e-7 1.6518e-5 1.5274e-4 3.5930e-4 1.1752e-3 3.3724e-3 8.8167e-3 4.6402e-12 3.6479e-11
Selective Training 25% 8.4251e-10 6.7845e-9 2.8456e-8 8.2103e-8 1.3825e-7 5.4620e-7 8.5305e-7 2.7764e-16 2.7764e-16
Selective Training 50% 2.4651e-14 1.1399e-13 5.4994e-13 2.7678e-12 1.4524e-11 7.9361e-11 4.5035e-10 2.0508e-24 1.8769e-22
Selective Training 75% 6.4234e-15 7.994e-14 2.9476e-13 1.1195e-12 4.3792e-12 1.7635e-11 7.3040e-11 1.6196e-25 1.0655e-24
Data Mixture 1:9 1.5110e-2 1.9836e-2 2.2986e-2 2.6625e-2 3.1173e-2 3.6624e-2 4.1384e-2 1.6196e-10 2.7675e-9
Data Mixture 1:3 6.7164e-3 1.0902e-2 1.4482e-2 1.9574e-2 2.7158e-2 3.2463e-2 3.7265e-2 6.9168e-21 2.2135e-19
Data Mixture 1:1 5.8732e-3 7.4283e-3 8.4372e-3 9.5638e-3 1.7128e-2 2.5756e-2 2.6832e-2 9.4685e-25 2.0132e-23
Data Mixture 3:1 5.1747e-4 9.2419e-4 1.2042e-3 1.4653e-3 1.6482e-3 1.9252e-3 2.3728e-3 9.6799e-48 2.4526e-45
IV-I4 Sensitivity Analysis of Detection Thresholds in Adversarial Attack Scenarios
Leveraging statistical tests under consistent settings, our method demonstrates exceptional effectiveness and robustness across various adversarial scenarios in the above discussion. Moreover, we evaluate the sensitivity of ModelShield to different thresholds across various adversarial scenarios. Table XII presents the $p$-values obtained from hypothesis testing in editing attack, selective training and mixture of unwatermarked data scenarios using our detection methods: both the rapid detection under different threshold settings and the detailed detection.
Text Editing Attacks: Both rapid and detailed watermark detection methods demonstrated high effectiveness and sensitivity across various detection thresholds. On the WILD dataset with Llama2, all rapid detection $p$-values were below 1e-10, while detailed detection $p$-values remained consistently below 1e-18.
Selective Training on Watermarked Data: Both rapid and detailed detection methods sensitively identified the presence of watermarks in various proportions and thresholds, revealing a distinct watermark pattern compared to the legitimate model. The $p$-values from rapid detection were all below 1e-2, and those from detailed detection were consistently below 1e-10.
Mixed Watermarked and Non-Watermarked Data: We tested the effectiveness of rapid and detailed inspections with mixed ratios of 1:9, 1:3, 1:1, and 3:1. Using Llama2 as the base model on the WILD dataset, the $p$-values from the rapid detection were consistently below 0.05, while the $p$-values from the detailed KS test were all below 1e-8 across various detection thresholds.
Under various attack scenarios, our rapid detection method remains effective even when the selected threshold fluctuates by up to 40%. ModelShield reliably identifies IP infringement across a range of thresholds, with most settings producing $p$-values below 0.001, demonstrating its robustness and insensitivity to threshold variations. Besides, Unlike the t-test used in rapid detection, the KS test employed in detailed verification does not rely on a predefined threshold. Instead, the $p$-value is calculated directly and compared to 0.05 for decision-making. The detailed verification results consistently demonstrate significant detection with $p<10^{-9}$ across all scenarios. This highlights ModelShield’s reliability and stability in identifying IP-infringing imitation models, even when subjected to various attack scenarios.
IV-J LLM Generation Visualization
@@
As illustrated in Fig 11, we analyze the BertScore [64] comparisons between the imitation model’s outputs and the original foundational model’s responses to identical queries. The imitation model consistently shows high sentence watermark scores across different scenarios, proving the watermark transfers effectively and is reliably detectable. Conversely, outputs from base models (represented in blue) show generally low sentence watermark scores, proving that the watermark is not prone to false positives.
@@ V Conclusion
In this paper, we propose ModelShield, a plug-and-play watermarking method for language models designed to combat model extraction attacks. Unlike existing IP protection watermarking methods that modify the model’s output or distribution with heuristic rules, ModelShield uses system prompt guidance to minimize distortion seamlessly. Leveraging the self-reminding capabilities of LLMs, ModelShield integrates user queries with a self-generated watermark system prompt, subtly embedding watermark words into outputs without altering the model’s internal probabilities or outputs. This approach maintains text quality while ensuring watermark effectiveness and robustness. Extensive experiments validate ModelShield’s efficacy, showing remarkable detection sensitivity even under interference. Its imperceptibility and model-agnostic nature make it practical for real-world applications. We hope ModelShield offers fresh insights for copyright protection of large language models. It provides a viable and efficient strategy for model owners to safeguard their IP rights at a low cost, fostering fairness and innovation in the LMaaS market.